Sign in
Please select an account to continue using cracku.in
↓ →
There are two bags, one of which contains 3 black and 4 white balls, while the other contains 4 black and 3 white balls. A dice is cast. If the face 1 or 3 turns up, a ball is taken from the first bag and if any other face turns up a ball is chosen from the second bag. The probability of choosing a black ball is
Bag1 = 3 back, 4 white balls
Bag2 = 4 black, 3 white balls
Probability of selecting bag 1 = 2/6 = 1/3
Probability of selecting bag 2 = 1 - 1/3 = 2/3
Probability of choosing a black ball from bag 1 = 3/7
Probability of choosing a black ball from bag 2 = 4/7
Required probability = 1/3 * 3/7 + 2/3 * 4/7
= $$\frac{11}{21}$$
Was this solution helpful?
Create a FREE account and get:
Book Free CAT Mentorship
Get personalized CAT strategy from a 99%iler
500+ students mentored
OTP Verification
Enter the 6-digit code sent to your phone
Booking Summary
Enter OTP
Didn't receive the OTP?
Start your IIM journey with the right preparation and crack CAT 2026.
Educational materials for CAT preparation